# Solving Empirical Bayes via Transformers

### v1

@@ 4 Understanding transformers
In this section, we try to gain an intuition on how transformers work in our setting. We achieve this from two angles. First, we establish some theoretical results on the expressibility of transformers in solving empirical Bayes tasks. Second, we use linear probes to study the prediction mechanism of the transformers.
@@ 4.1 Expressibility of Transformers
We discuss the feasibility of using transformers to solve the empirical Bayes prediction task. Indeed, the study of the universal approximation theorem has been done on multilayer perceptron, c.f. [Aug24], with some variations like bounded weights [GI18] and width [KL20, PYLS20]. More recently, universal learnability of transformers has been established, first in [YBR+19], which shows that 2 heads, each of size 1, and 4 hidden dimensions are all we need. [FdHP24] further characterizes universal learnability in terms of in-context learning.
@@
Set a positive integer $d$ and a positive real number $M$. Then for any $\epsilon>0$, there exists a transformer architecture with one encoding layer, skip connection, and embedding dimension $d+1$ that learns the clipped Robbins estimator $\hat{\theta}_{\mathsf{Rob},d,M}$ up to a precision $\epsilon$.
Similarly, we may show that transformers can approximate NPMLE up to an arbitrary input value and precision.
@@
Let $M>0$, and denote the NPMLE estimator $\hat{\theta}_{\mathsf{NPMLE},M}$, the NPMLE estimator chosen among $\mathcal{P}([0,M])$. For each integer $d>0$ consider the following modified NPMLE function:
@@
then for any $\epsilon>0$ there exists a transformer network that can approximate $\theta_{\mathsf{NPMLE},d}$ uniformly up to $\epsilon$-precision.
Full proofs are deferred to Appendix B and we only give a sketch for now. For Robbins approximation, we create an encoding mechanism that encodes $\frac{N(X_{i})}{N(X_{i})+(X_{i}+1)N(X_{i}+1)}$ at position $i$ among $1,\cdots,n$ and use a decoder to approximate the function $x\to\min\{\frac{1}{x}-1,M\}$. For NPMLE approximation, we pass in the Sigmoid of the integer inputs as embedding, and show that $\hat{\theta}_{\mathsf{NPMLE}}$ can be continually extended, with sigmoid-transformed empirical distribution as arguments. For the encoding part, we provide a pseudocode in Appendix C that closely follows PyTorch’s implementation.
To illustrate the significance of both of these theorems, we demonstrate that transformers can learn an empirical Bayes prediction task to an arbitrarily low regret.
@@
For any $\epsilon>0$, there exists an integer $N$ and a transformer network $\Gamma$ such that for all $n\geq N$, the minimax regret of $\Gamma(X_{1},\cdots,X_{n})$ on prior $\pi\in\mathcal{P}([0,\theta_{\max}])$ satisfies
@@ 4.2 How do transformers learn?
We study the mechanisms by which transformers learn via linear probe [AB18]. To this end, we take the representation of each layer of our pretrained transformers, and train a decoder that comprises a layer normalization operation, linear layer, and GeLU activation. This decoder is then trained with the following labels: frequency $N(x)$ within a sequence, and posterior density $f_{\hat{\pi}}(x)$ estimated by the NPMLE. The aim is to study whether our transformers function like the Robbins or NPMLE. In the plot in Fig. 1, we answer this as negative, showing that our transformers are not merely learning about these features, but instead learning what the Bayes estimator $\hat{\theta}_{\pi}$ is.
Figure 1: $R^{2}$ score of linear probe results against $N(x)$,$f_{\pi}(x)$ and $x$ for T18 (the plots for T24 appear similar). We see that while $x$ itself is easily recoverable from any layer, “knowledge” about the former two quantities appears to decrease with depth.
@@ 5 Synthetic Experiments
We now evaluate our trained transformers on the following: How well do they generalize? This can be done by evaluating on the following: sequence lengths other than the ones we have trained on, unseen priors, and unknown bound $\theta_{\max}$. We also compare against the classical algorithms introduced in Section 3 to demonstrate the superiority of these transformers by showing the average regret; most other details are deferred to Appendix D.1. We also investigate the inference time to show its advantage over NPMLE.
@@ 5.1 Ability to Generalize
Adaptibility to various sequence lengths. In this experiment, we evaluate the ability of transformers to adapt to different sequence lengths, both fewer than and more than what is trained. To do so, we evaluate them on 4096 neural prior-on-priors (which is part of the training distribution), but on various sequence length $n$: 128, 256, 512, 1024, and 2048. For each such prior, we generate 192 batches for evaluation. We report the average regret over the 4096 priors in Fig. 2.
Figure 2: Regret vs sequence length (neural prior). The regret decreases for both transformers as the sequence length increases, showing that they do have the ability to generalize. We nevertheless note that NPMLE has a better generalization ability, as shown by the regret at sequence length 2048 as compared to smaller sequences. In comparison, the average regret for ERM monotone is 11.20, 8.19, 5.58, 3.66, and 2.36 for the various sequence lengths, while the average regret for MLE and GS stays constant at 14.816 and 14.658, respectively.
@@
Robustness against unseen priors. Our transformers are trained on a mixture of neural and Dirichlet priors. Here, we consider their performances on the worst case prior in $\mathcal{P}([0,50])$ as mentioned in Section 3.1 and further explained in Appendix A.1. The numbers of batches we use in this prior are 786k (for sequence lengths $n=128,256,512$), 393k (for $n=1024$), and 197k (for $n=2048$). We also consider another unseen prior-on-prior: the multinomial prior supported on $[0,50]$ with fixed, evenly split grids and weights distributed as Dirichlet distribution, using sequence lengths 512, 1024, and 2048, using 192 batches for each of the 4096 priors we evaluate on. We report the estimated regret in Fig. 3 and Fig. 4 to show that transformers produce regret comparable to the strongest alternative (NPMLE).
Figure 3: Regret of various transformers on worst prior compared to NPMLE. Again, transformers show the ability to generalize to longer sequence lengths, although for longer sequences NPMLE generalizes better. Note that this is already better than ERM-monotone’s regret at 12.79, 9.80, 6.99, 4.81, and 3.26 across the 5 sequence lengths, while MLE’s regret stays at 11.73.
Figure 4: Average regret of various transformers on multinomial priors compared to NPMLE. Here, transformers show $\geq 2$ times improvement over NPMLE even on sequence lengths it never trained on. In comparison, ERM-monotone’s regret across sequence lengths are at 5.17, 3.69, 2.57, while MLE’s and GS’s regrets stay at 5.87 and 5.63, respectively.
Training under randomized $\theta_{\max}$. In another experiment, we investigate the effect of mismatched $\theta_{\max}$. on the performance of transformers without knowledge of $\theta_{\max}$. Specifically, we train two sets of transformers, one as reported in Table 1, the other set (T18r, T24r) with the same parameters but with $\theta_{\max}$ randomized according to the following mixture:
$\theta_{\max}\sim\frac{3}{4}\mathcal{N}(0,50,10^{2})+\frac{1}{8}\text{Exp}(50)+\frac{1}{8}\text{Cauchy}(50,10)$
and clamped at $[10,150]$. Then, for the two sets of transformers, we evaluate them on 4096 neural prior-on-priors, using the default sequence length = 512 and 192 batches for each prior. We report the distribution of regrets in Fig. 5 which demonstrates that transformers trained with randomized $\theta_{\max}$ see a small deterioration in regret, but nonetheless still outperform NPMLE in regret minimization.
Figure 5: Comparison of regrets when $\theta_{\max}$ is trained randomly. The mean regret increases by 18.5% and 23.4% for the transformers with 18 and 24 layers, respectively, when compared against transformers that learn the true $\theta_{\max}$ during training, but still outperforms NPMLE. All comparisons resulted in a significant $t$-score ($p<$1e-100).
5.2 Inference Time Comparison
We evaluate their inference time of various estimators over 4096 neural prior-on-priors, where for each prior we consider the time needed to estimate the hidden parameter of 192 batches and sequence length 128, 256, 512, 1024, and 2048. Each program is given 2 Nvidia Volta V100 GPUs and 40 CPUs for computation. The results are tabulated at Fig. 6, where we see that the transformers’ runtime is comparable to that of ERM’s.
Figure 6: Time vs sequence length, showing that the inference time of transformers is comparable with that of ERM monotone. We nevertheless qualify this finding by noting that these running times seem to scale superlinearly with sequence length. For comparison, the running time of NPMLE for sequence lengths 128, 256, 512, 1024, and 2048 are 41.69, 67.70, 109.81, 175.72, and 289.79 seconds, respectively, which indicates that transformers are 2 orders of magnitudes faster than NPMLE.
@@ 6 Real Data Experiments
In this section, we answer the following question: Can our transformers that are pre-trained on synthetic data perform well on real datasets without re-training on any part of the real datasets?
To do so, we consider the following experimental setup: Given an integer-valued attribute, let $X$ be the count of the attribute in the initial section we observe, and $Y$ be the count of a similar attribute in the remaining section that we should predict. We assume that given a horizon length (duration, sentence length, etc) $n_{X}$ and $n_{Y}$ of the two sections, there exist hidden parameters $\theta_{i}$ such that $X_{i}\sim\text{Poi}(n_{X}\theta_{i})$ and $Y_{i}\sim\text{Poi}(n_{Y}\theta_{i})$, independently (for convenience we will scale $\theta_{i}$ such that $n_{X}=1$). Our goal is to predict $\hat{Y}=n_{Y}\hat{\theta}(X)$ using empirical Bayes methods. We will focus on the following two types of datasets: sports and word frequency. Below, we describe the types of datasets that we would study. Throughout this section, we name $(X,Y)$ as the input and label sets, respectively.
@@ 6.1 Sports datasets
NHL dataset. We proceed in the same spirit as [JPW22], Section 5.2, and study the data on the total number of goals scored by each player in the National Hockey League for 29 years: from the 1989-1990 season to the 2018-2019 season (2004-2005 season was canceled). The data is obtained from [Hoc00], and we focus on the skaters’ statistics. Here, given the number of goals a player scored in season $j$, we wish to predict the same for season $j+1$ (thus the input and label sets are the number of goals a player scored in consecutive seasons, and $n_{Y}=1$). We study the prediction results when fitting all players at once, as well as fitting only positions of interest (defender, center, and winger).
MLB dataset. The dataset is publicly available at [Ret96], and can be processed by [Est18]. Here, we study the hitting count of each player in batting and pitching players from 1990 to 2017. Unlike the between-season prediction as we did for the NHL dataset, we do in-season prediction. That is, we take $X$ as the number of goals scored by a player in the beginning portion of the season, and $Y$ in the rest of the season. For batting and pitching players (which we fit separately), we use $X$ as the goals in the first $\frac{1}{5}$ and first $\frac{1}{6}$ of the season (i.e. $n_{Y}=4,5$), respectively.
@@ 6.2 Word frequency datasets
In this setting, we model the alphabet of tokens as $M$ categorical objects $A=\{A_{1},\cdots,A_{M}\}$. Given $n$ samples from these objects, and denote $(X_{1},\cdots,X_{M})$ the frequency of the samples. Suppose we are to estimate the frequencies $(Y_{1},\cdots,Y_{M})$ of an unseen section of length $t$ (here $t$ known). We model as follows: consider $p_{1},\cdots,p_{M}$ as the “inherent” probability distribution over $M$ (or proportion in a population), so $\sum_{i=1}^{M}p_{i}=1$. Now the frequency $X_{i}\sim\text{Binom}(n,p_{i})$, which we may instead approximate as $X_{i}\sim\text{Poi}(np_{i})$. Thus we may use empirical Bayes method to estimate $\hat{\theta}_{i}=n\hat{p}_{i}$ based on the frequencies $X_{1},\cdots,X_{M}$, and then predict $\hat{Y}_{i}=\frac{t}{n}\hat{\theta}_{i}$.
@@ 6.3 Evaluation Methods
Significance of improvement. We will also consider whether one improvement is significant by performing paired $t$-test on the RMSE of transformers against the baselines.
@@
We tabulate our findings in Table 2 and Table 3. In addition, we also show a few violin plots in Fig. 7, Fig. 8, Fig. 9 for NHL, MLB batting, and Bookcorpus to supplement Table 2 (with Robbins removed due to its wide variance). From Table 3, we conclude a nontrivial improvement of the transformers over the classical methods in most of the datasets. A more detailed comparison (e.g. the ELO rating of estimators’ RMSE and the MAE metric), is shown in Appendix D.2.
Figure 7: Violin plots of RMSE ratio achieved by multiple estimators over MLE on NHL.
Figure 8: Violin plots of RMSE ratio achieved by multiple estimators over MLE on MLB batting.
Figure 9: Violin plots of RMSE ratio achieved by multiple estimators over MLE on BookCorpusOpen.
Table 2: 95% confidence interval of the percentage improvement of RMSE by each algorithm over MLE.
Dataset Robbins ERM NPMLE T18 T24
NHL -30.55 $\pm$ 6.55 1.46 $\pm$ 0.65 3.24 $\pm$ 0.92 3.51 $\pm$ 1.01 3.46 $\pm$ 1.00
NHL (defender) -19.54 $\pm$ 6.35 3.19 $\pm$ 1.32 6.48 $\pm$ 1.63 7.25 $\pm$ 1.88 7.41 $\pm$ 1.86
NHL (center) -49.89 $\pm$ 10.36 0.38 $\pm$ 0.82 3.44 $\pm$ 0.94 4.12 $\pm$ 1.14 4.06 $\pm$ 1.07
NHL (winger) -42.63 $\pm$ 7.58 0.76 $\pm$ 0.69 3.06 $\pm$ 0.87 3.39 $\pm$ 1.03 3.38 $\pm$ 1.01
MLB (batting) -32.80 $\pm$ 5.67 2.50 $\pm$ 0.36 4.30 $\pm$ 0.41 4.45 $\pm$ 0.37 4.58 $\pm$ 0.39
MLB (pitching) -21.71 $\pm$ 2.45 2.51 $\pm$ 0.31 4.70 $\pm$ 0.41 4.89 $\pm$ 0.42 4.95 $\pm$ 0.38
BookCorpusOpen -4.58 $\pm$ 0.43 9.38 $\pm$ 0.10 10.82 $\pm$ 0.11 10.38 $\pm$ 0.18 11.43 $\pm$ 0.17
[…]
T18 T24
Dataset MLE Robbins ERM NPMLE MLE Robbins ERM NPMLE
NHL 1.44e-06 2.51e-11 8.42e-06 0.120 1.57e-06 3.14e-11 8.96e-06 0.150
NHL (defender) 2.15e-08 9.76e-11 4.71e-08 8.72e-05 1.48e-08 9.67e-11 3.54e-08 4.77e-05
NHL (center) 1.14e-06 1.33e-11 6.01e-07 2.98e-03 6.38e-07 1.73e-11 5.61e-07 1.65e-03
NHL (winger) 1.76e-06 8.01e-13 1.33e-05 0.153 1.54e-06 8.29e-13 1.16e-05 0.152
MLB batting 1.39e-21 6.51e-14 1.30e-11 1.08e-03 8.49e-22 6.21e-14 1.93e-12 2.11e-08
MLB pitching 9.62e-20 1.21e-16 1.49e-12 2.34e-03 5.95e-21 1.13e-16 2.10e-13 4.78e-06
BookCorpusOpen $<$ 1e-100 $<$ 1e-100 0.0104 1 - 1.19e-05 $<$ 1e-100 $<$ 1e-100 6.88e-15 0.133
7 Conclusion and Future Work

### v2

@@ 4 Understanding transformers
In this section, we try to gain an intuition on how transformers work in the empirical Bayes setting. We achieve this from two angles. First, we establish some theoretical results on the expressibility of transformers in solving empirical Bayes tasks. Second, we use linear probes to study the prediction mechanism of the transformers.
@@ 4.1 Expressibility of transformers
We discuss the feasibility of using transformers to solve the empirical Bayes prediction task. Universal learnability of transformers is first established in [YBR+19], and further characterized in [FdHP24].
@@
Set a positive integer $d$ and a positive real number $M$. Then for any $\epsilon>0$, there exists a transformer that learns the clipped Robbins estimator $\hat{\theta}_{\mathsf{Rob},d,M}$ up to a precision $\epsilon$, In addition, this transformer has embedding dimension $d+1$ at encoding stage, and a two-layered feedforward network with $O((1+M)^{2}\epsilon^{-1})$ hidden neurons at the hidden layer at decoding stage.
We also show that transformers can approximate NPMLE up to an arbitrary input value and precision, although we will omit quantitative bounds on this transformer due to challenges stated in [FdHP24, Remark 3].
@@
Let $M>0$, and denote the NPMLE estimator $\hat{\theta}_{\mathsf{NPMLE},M}$, the NPMLE estimator chosen among $\mathcal{P}([0,M])$. For each integer $d>0$, consider the following modified NPMLE function:
@@
then for any $\epsilon>0$ there exists a transformer that can approximate $\theta_{\mathsf{NPMLE},d}$ uniformly up to $\epsilon$-precision.
Full proofs are deferred to Appendix B, and we only give a sketch for now. For Robbins approximation, we use softmax attention to create an encoding mechanism that encodes $\frac{N(X_{i})}{N(X_{i})+(X_{i}+1)N(X_{i}+1)}$ at position $i$ among $1,\cdots,n$ and use a decoder to approximate the function $x\to\min\{\frac{1}{x}-1,M\}$. For NPMLE approximation, we pass in the Sigmoid of the integer inputs as embedding, and show that $\hat{\theta}_{\mathsf{NPMLE}}$ can be continuously extended, with sigmoid-transformed empirical distribution as arguments. Then by [FdHP24, Theorem 1], there exists a transformer that can learn NPMLE to an arbitrary precision. For the encoding part, we provide a pseudocode in Appendix C that closely follows PyTorch’s implementation.
To illustrate the significance of both of these theorems, we demonstrate that transformers can learn an empirical Bayes prediction task to an arbitrarily low regret, although the quantitative bounds for the transformer size only apply for Robbins estimation.
@@
Let $\theta_{\max}>0$ be fixed. For any $\epsilon>0$, there exists an integer $N=O_{\theta_{\max}}(\epsilon^{-1}\log^{2}(\epsilon^{-1})))$ and a transformer $\Gamma$ with $O_{\theta_{\max}}(\epsilon^{-1})$ parameters such that for all $n\geq N$, the minimax regret of $\Gamma(X_{1},\cdots,X_{n})$ on prior $\pi\in\mathcal{P}([0,\theta_{\max}])$ satisfies
@@ 4.2 How do transformers learn?
We study the mechanisms by which transformers learn via linear probe [AB18]. To this end, we take the representation of each layer of our pretrained transformers, and train a decoder that comprises a layer normalization operation, a linear layer, and GeLU activation. This decoder is then trained with the following labels: frequency $N(x)$ within a sequence, and posterior density $f_{\hat{\pi}}(x)$ estimated by the NPMLE. The aim is to study whether our transformers function like the Robbins or NPMLE. In the plot in Fig. 1, we answer this as negative, showing that our transformers are not merely learning about these features, but instead learning what the Bayes estimator $\hat{\theta}_{\pi}$ is.
(a) Neural prior
(b) Multinomial prior
(c) Multinomial prior: PMF($\theta$)
Figure 1: (a), (b): $R^{2}$ score of linear probe result against $N(x),f_{\hat{\pi}}(x)$ and $x$ for T24r. We see that while $x$ itself is easily recoverable from any layer, “knowledge” about the former two quantities appears to either decrease (in (a)) or plateau (in (b)) with depth. (c) In the multinomial prior case, T24r does not seem to use any information on the atom weight $\text{PMF}_{\pi}(\theta)$.
@@ 5 Synthetic experiments
We now evaluate our trained transformers on their ability to generalize to sequence lengths and priors unseen during the training stage. We also compare against the classical algorithms introduced in Section 3 to demonstrate the superiority of these transformers by showing the average regret; most other details are deferred to Appendix E. We also investigate the inference time to show its advantage over the classical algorithms.
@@ 5.1 Ability to generalize
Our transformer T24r is trained on a mixture of neural and Dirichlet priors, with randomized $\theta_{\max}$, and of sequence length 512. We will now describe the aspects we evaluate our trained transformer T24r on, and the corresponding experiments to do so. We will also evaluate L24r to investigate the efficacy of linear attention transformers.
Adaptability to various sequence lengths. We evaluate the ability of transformers to adapt to different sequence lengths, both fewer than and more than what is trained. For each of the experiments, we evaluate our transformers and the baselines on sequence lengths 128, 256, 512, 1024, and 2048.
@@
Robustness against unseen priors. We evaluate our transformers on the following two priors unseen during training: (a) worst case prior in $\mathcal{P}([0,50])$ as mentioned in Section 3.1 and further explained in Appendix A.1; (b) multinomial prior-on-priors supported on $[0,50]$ with fixed, evenly split grids and weights following Dirichlet distribution. For the worst case prior, the numbers of batches we evaluate on for this prior are 786k (for sequence lengths $n=128,256,512$), 393k (for $n=1024$), and 197k (for $n=2048$). For multinomial prior-on-priors, we evaluate on 4096 priors, with the number of batches for each prior set as 4000 ($n=128,256,512$) and 2000 ($n=1024,2048$), except for NPMLE where we use 192 batches for each prior given the significant computational resource needed.
Robustness under unknown $\theta_{\max}$. We investigate the following: how much can we gain with perfect information on $\theta_{\max}$ during training? Here, we evaluate our transformer T24r on 4096 neural-generated prior-on-priors with $\theta_{\max}=50$, and compare with another transformer T24f with the same parameters, with training prior still neural-Dirichlet mixture (as described in Section 3.3) but fixed $\theta_{\max}=50$. The numbers of batches we evaluate on for each of the 4096 priors are 4000 (for sequence lengths $n=128,256,512$) and 2000 (for sequence lengths $n=1024,2048$), except for NPMLE where we use 192 batches per prior. In the plots in Fig. 2 we see that removing information on $\theta_{\max}$ (T24r vs T24f) during training incurs around 40% extra regret.
Analysis of results. We show our results in Fig. 2, which compares NPMLE, T24r, L24r, and T24f (when applicable). Note that we also run against other baselines (ERM-monotone, GS, MLE, Robbins), but their regrets are too big to be shown in the plots. There are situations where L24r outperforms T24r, showing that linear transformers perform almost as well. However, T24r appears to show a better length generalization property.
(a) Worst case prior
(b) Multinomial prior-on-priors
(c) Neural prior-on-priors
Figure 2: Average regret of NPMLE vs transformers on various priors in $\mathcal{P}([0,50])$ and on various sequence lengths. The regrets of T24r and L24r decrease with sequence length on all priors and beats NPMLE on most instances. Red vertical lines at 512 denotes the sequence length that the transformers are trained on. (a). T24r outperforms L24r at longer sequence length. ERM-monotone’s, MLE’s, and Robbin’s regrets are 3.26, 11.73, and 85.68 even at sequence length 2048. (b). L24r outperforms T24r throughout, although the gap narrows at $n=2048$. ERM-monotone’s, MLE’s, GS’s, and Robbin’s regrets are 2.57, 5.87, 5.63, and 64.43 even at sequence length 2048. (c). NPMLE generalizes better at longer sequence lengths. At $n=4096$ (not shown) NPMLE beats the best performing T24f (regret 0.104 vs 0.153). T24r has around 40% extra regret compared to T24f throughout; L24r has 10% extra regret compared to T24r at $n=2048$. ERM-monotone’s, MLE’s, GS’s and Robbin’s regrets are 2.36, 14.82, 14.66, and 69.40 even at sequence length 2048.
5.2 Inference time analysis
Figure 3: Average time (in seconds) per batch vs sequence length, showing that the inference time of T24r is comparable with that of ERM monotone, and 100x faster than NPMLE. Also shown is L24r which scales better at $n=2048$.
The inference time of T24r in terms of sequence length $n$, number of layers $L$, and embedding dimension is $O(Ln^{2}d)$. In contrast, the time complexity of NPMLE (the most competitive classical algorithm) is difficult to obtain: the algorithms were either based on expectation-maximization (which does not have provable guarantees of convergence) or convex optimization (where the rigorous study is also in its infancy [PS25]).
To compare T24r against classical EB algorithms, we run our estimators over 4096 neural prior-on-priors, where for each prior we consider the time needed to estimate the hidden parameter of 192 batches and sequence lengths $n=$128, 256, 512, 1024, and 2048. Each program (transformer and classical baselines) is given 2 Nvidia Volta V100 GPUs and 40 CPUs for computation (340 GB total RAM). The results are shown at Fig. 3, where we see that T24r is faster than ERM-monotone on GPU and 100x faster than NPMLE. We attribute this to that T24r can better leverage the compute power of GPUs compared to classical algorithms. For interest, we will also attach the running time of L24r, which is linear in complexity and runs 10x faster than T24r when $n=2048$.
@@ 6 Real data experiments
In this section, we answer the following question: Can our transformers that are pre-trained on synthetic data perform well on real datasets without re-training on any part of the real datasets, and without any prompt in the form of input-label pairs?
To do so, we consider the following experimental setup: Given an integer-valued attribute, let $X$ be the count of the attribute in the initial section we observe, and $Y$ be the count of a similar attribute in the remaining section that we should predict. We assume that given a horizon length (duration, sentence length, etc) $n_{X}$ and $n_{Y}$ of the two sections, there exist hidden parameters $\theta_{i}$ such that $X_{i}\sim\text{Poi}(n_{X}\theta_{i})$ and $Y_{i}\sim\text{Poi}(n_{Y}\theta_{i})$, independently (for convenience we will scale $\theta_{i}$ such that $n_{X}=1$). Our goal is to predict $\hat{Y}=n_{Y}\hat{\theta}(X)$ using empirical Bayes methods. We will focus on the following two types of datasets: sports and word frequency. These are some of the main focuses of EB literature [Efr12, Tables 1.1, 11.6]. In the following, we describe the types of datasets that we would study. Throughout this section, we name $(X,Y)$ as the input and label sets, respectively.
We will compare against the classical EB methods. While deep-learning Bayesian models like Generative Bayesian Computation [PRS24], TabPFN [HMP+25], Simformer [GDW+24], Amortized Conditioning Engine [CLH+25] appear to solve empirical Bayes, this is not the case, which we will explain more thoroughly in Appendix F.1.
@@ 6.1 Sports datasets
NHL dataset. We proceed in the same spirit as [JPW22, Section 5.2], and study the data on the total number of goals scored by each player in the National Hockey League for 29 years: from the 1989-1990 season to the 2018-2019 season (2004-2005 season was canceled). The data is obtained from [Hoc00], and we focus on the skaters’ statistics. For each season $j$ we predict the goals scored by a player in season $j+1$ based on their score in season $j$. (thus the input and label sets are the number of goals a player scored in consecutive seasons, and $n_{Y}=1$). We study the prediction results when fitting all players at once, as well as fitting only positions of interest (defender, center, and winger).
MLB dataset. The dataset is publicly available at [Ret96], and can be processed by [Est18]. Here, we study the hitting count of each player in batting and pitching players from 1990 to 2017. Unlike the between-season prediction as we did for the NHL dataset, we do in-season prediction. That is, we take $X$ as the number of goals scored by a player in the beginning portion of the season, and $Y$ in the rest of the season. For both batting and pitching players (which we fit separately), we use $X$ as the goals in the first half of the season (i.e. $n_{Y}=1$).
@@ 6.2 Word frequency datasets
In this setting, we model the alphabet of tokens as $M$ categorical objects $A=\{A_{1},\cdots,A_{M}\}$. Given $n$ samples from these objects, we denote $(X_{1},\cdots,X_{M})$ the frequency of the samples. We are to estimate the frequencies $(Y_{1},\cdots,Y_{M})$ of an unseen section of length $t$ (here $t$ known). We model as follows: consider $p_{1},\cdots,p_{M}$ as the “inherent” probability distribution over $M$ (or proportion in a population), so $\sum_{i=1}^{M}p_{i}=1$. Now the frequency $X_{i}\sim\text{Binom}(n,p_{i})$, which we may instead approximate as $X_{i}\sim\text{Poi}(np_{i})$. Thus we may use empirical Bayes method to estimate $\hat{\theta}_{i}=n\hat{p}_{i}$ based on the frequencies $X_{1},\cdots,X_{M}$, and then predict $\hat{Y}_{i}=\frac{t}{n}\hat{\theta}_{i}$.
@@ 6.3 Evaluation methods
Significance of improvement. We consider whether one improvement is significant by performing paired $t$-test on the RMSE of transformers against the baselines and report the $p$-value.
@@
We will only show percentage improvement over MLE in Table 1; $t$-tests and Plackett-Luce ranking are deferred to Appendix F. We also show violin plots in Fig. 4 to supplement our findings (with Robbins removed due to its wide variance). A more detailed comparison, including the use of the MAE metric, is shown in Appendix F. These experiments seem to support that transformers, including the much faster L24r, can outperform the classical EB estimators in real-life settings.
(a) NHL Hockey
(b) MLB (Batting)
(c) MLB (Batting)
(d) BookCorpusOpen
Figure 4: Violin plots of RMSE ratio of ERM-monotone (blue), NPMLE (orange), T24r (green), and L24r (red) over MLE over various datasets. The horizontal line at 1.0 indicates the threshold of MLE’s RMSE. Robbins is not shown due to the wide variance. Both T24r and L24r show a general improvement over MLE, and in the case of the MLB dataset their advantage over NPMLE is visible.
Table 1: 95% confidence interval of the percentage improvement of RMSE over MLE.
Dataset Robbins ERM NPMLE T24r L24r
NHL (all) -30.55 $\pm$ 6.55 1.46 $\pm$ 0.65 3.21 $\pm$ 0.92 3.46 $\pm$ 0.88 3.96 $\pm$ 1.12
NHL (defender) -19.54 $\pm$ 6.35 3.19 $\pm$ 1.32 6.48 $\pm$ 1.63 6.91 $\pm$ 1.71 7.54 $\pm$ 2.04
NHL (center) -49.89 $\pm$ 10.36 0.38 $\pm$ 0.82 3.44 $\pm$ 0.94 4.02 $\pm$ 0.99 4.32 $\pm$ 1.22
NHL (winger) -42.63 $\pm$ 7.58 0.76 $\pm$ 0.69 3.06 $\pm$ 0.87 3.44 $\pm$ 0.89 3.76 $\pm$ 0.99
MLB (batting) -59.66 $\pm$ 5.88 0.23 $\pm$ 0.18 1.25 $\pm$ 0.18 1.50 $\pm$ 0.16 1.44 $\pm$ 0.17
MLB (pitching) -40.81 $\pm$ 3.16 0.09 $\pm$ 0.14 1.21 $\pm$ 0.19 1.42 $\pm$ 0.18 1.38 $\pm$ 0.17
BookCorpusOpen -4.58 $\pm$ 0.43 9.38 $\pm$ 0.10 10.82 $\pm$ 0.11 9.43 $\pm$ 0.12 11.23 $\pm$ 0.21
7 Conclusion, limitation and future work